$SPAD/src/input torus.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool torus.output
)set message test on
)set message auto off
)clear all
--S 1 of 3
f(x:SF):SF == x
--R
--R   Function declaration f : DoubleFloat -> DoubleFloat has been added
--R      to workspace.
--R                                                                   Type: Void
--E 1
--S 2 of 3
torus : TUBE := tubePlot(sin t,cos t,0,f,0..2*%pi,0.5::SF,12,"closed")
--R
--R
--RDaly Bug
--R   Although TubePlot is the name of a constructor, a full type must be
--R      specified in the context you have used it. Issue )show TubePlot
--R      for more information.
--E 2
--S 3 of 3
makeViewport3D(torus,"torus")$VIEW3D
--R
--R   There are 2 exposed and 0 unexposed library operations named
--R      makeViewport3D having 2 argument(s) but none was determined to be
--R      applicable. Use HyperDoc Browse, or issue
--R                         )display op makeViewport3D
--R      to learn more about the available operations. Perhaps
--R      package-calling the operation or using coercions on the arguments
--R      will allow you to apply the operation.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      makeViewport3D with argument type(s)
--R                                   Symbol
--R                                   String
--R

--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 3
)spool
)lisp (bye)

———-

References^{[1] nothing}

5